OpenStudy (anonymous):

help me pls: using substitution -2x+y=7 -x+3y=11

OpenStudy (anonymous):

muliply eq. 2 by -2 ===> -2x+y=7 , 2x-3y=-22 ===> sum both ==> -2y = -15 ===> y=15/2 ===> sybtitute in eq 1 ==> -2x+15/2=7 ==> x=1/4

OpenStudy (anonymous):

can you show me the solution pls?

OpenStudy (anonymous):

ok, I'll carry out that.

OpenStudy (anonymous):



OpenStudy (anonymous):

Just get one equation in terms of a variable and plug that equation in wherever you see that variable in the second equaiton

OpenStudy (anonymous):

-2x+y=7 so y=7+2x then we replace the y of second equation with 7+2x because it is equal to y. -x+3(7+2x)=11 ----> multiply 7+2x by 3 -x+21+6x=11 -----> add the x and transfer 21 to other side 5x = -10 ---------> devide both side by 5 x = -2 now substitute the value of x to any equation to get the value of y -2(-2)+y=7 ----> multiply 4+y = 7 -------> transfer 4 to other side y = 3 so x = -2 and y = 3 :-)

OpenStudy (anonymous):

hope i have helped :-)

OpenStudy (anonymous):

checked your work; seems good to me :)

OpenStudy (anonymous):

i get the same too